OpenStudy (anonymous):

PLEASE HELP!!!!! I GIVE MEDALS!!!!!!! The distance traveled by an object can be modeled by the equation d = ut + 0.5at2 …where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work.

OpenStudy (anonymous):

d = ut + 0.5at^2 0.5at^2 = d - ut a = (d - ut)/(2t^2) \[a = \frac{ d - ut }{ 2t ^{2} }\] those are your steps leading to the answer. assuming you meant to put (0.5at^2) as in t squared. not t * 2

OpenStudy (anonymous):

YES!!! Thank you so much!!!! :)

OpenStudy (anonymous):

you're welcome.

OpenStudy (anonymous):

\[d=\frac{a t^2}{2}+t u \]\[d-t u=\frac{a t^2}{2} \]\[\frac{d-t u}{\frac{t^2}{2}}=a \]\[\frac{2 (d-t u)}{t^2}=a \]

OpenStudy (anonymous):

Thank you too @robtobey :) I'd give you a medal too if I could :)

OpenStudy (anonymous):

np